OpenStudy (anonymous):

Solve using the square root method. 9x2 - 25 = 0 A. {-5/3, 5/3} B. {-3/5, 3/5} C. {-3, 5} D. {-5, 3}

OpenStudy (anonymous):

i got B.

OpenStudy (hoblos):

a^2 - b^2 =)a-b)(a+b) 9x2 - 25 = 0 (3x-5)(3x+5) =0 x=5/3 or x=-5/3

OpenStudy (hoblos):

A

OpenStudy (anonymous):

I had it backwards i think :/